student174:

A shop recorded total sales of $2,000 on Monday. On Tuesday, its sales fell by 10%. On Wednesday, sales fell by another 20% compared with Tuesday. From Wednesday to Thursday, sales increased by an amount equal to 25% of Monday’s total sales. What is the net change, in dollars, in total sales from Monday to Thursday? The net change, in dollars, in total sales was

student174:

Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar.

Hoodmemes:

Well this is simple u wanna make the percentage a decimal \[10\div100=0.1\] \[0.1\times2000=200\] Since we already know they recorded the total sales of $2,000 on Monday. \[2000-200=1,800\] So the sales decreased at 200 dollars And they had 1,800 sales on Tuesday Now Wednesday \[20\div100=0.2\] Now multiply that by the total sales on Tuesday \[0.2\times1800=360\] So there was a 360 decrease on Wednesday \[1800-360=1440\] So they made 1440 sales on Wednesday Since it says there was a "sales increased by an amount equal to 25% of Monday’s total sales." \[25\div100=0.25\] \[0.25\times2000=500\] So there was a 500 dollar increase \[1440+500=1940\] So since we started with 2,000 on monday and ended up with 1940 on thursday to find the net change we must subtract \[2000 - 1940=60\] So the Net Change was by 60 dollars (Woops sorry this took a while)